OpenStudy (anonymous):

square root of -20 times square root of -27 ...no decimals

OpenStudy (anonymous):

sqrt(20)i*sqrt(27)i =sqrt(540)

OpenStudy (anonymous):

=3sqrt(60)

OpenStudy (kinggeorge):

Perhaps an easier way to see it...\[\sqrt{-20} \cdot \sqrt{-27} = \sqrt{(-20)(-27)}=\sqrt{540}=3\sqrt{60}\]

OpenStudy (anonymous):

would that simplify to 6 square root 15 ?

OpenStudy (kinggeorge):

Yes.